# mO Haoyu[#1]

### v1

@@ 3 The Challenge of Textural Separability
In the previous section, we demonstrated that SAM struggles with segmenting non-conventional shapes, in particular, dense tree-like structures, which we hypothesize is due to the model confusing the dense structure as the texture of a non-treelike, more regular shape, rather than a shape itself (see e.g. Fig. 1 left). Similar to this behavior is that even for objects with simpler shapes, SAM can still be confused if the object’s texture is even somewhat similar to its surroundings, which we will study in this section.
@@ 3.1 Measuring Textural Separability
We will define the textural contrast or separability between some object mask and its surroundings by how easily their textures can be distinguished from one another. Motivated by findings that early layers of classification-pretrained CNNs primarily capture low-level features involving edges and textures [44], we will characterize an image’s textures using the first convolutional layer of a ResNet-18 [19] pretrained on ImageNet [9]. Denoted by $f_{1}:\mathbb{R}^{C\times H\times W}\rightarrow\mathbb{R}^{C^{\prime}\times H^{\prime}\times W^{\prime}}$, this outputs a textural feature map corresponding to all $7\times 7$ windows in the image.
We then measure the textural separability of an object according to if a simple classifier $g$ can be trained to discriminate between (a) the activations for the pixels of the object foreground and (b) the activations right outside of the object boundary [^4] This procedure is somewhat similar to the probing of hidden activation concepts [2, 26], although these works detected concepts (such as textures) at the image level, not at the single activation level as we do here.. We define this explicitly in Algorithm 3, where $\mathrm{Dilate}$ and $\mathrm{Disk}$ refer to scikit-image.morphology functions [41]. We use simple logistic regression for $g$ with inverse regularization parameter $C=2$ [31]. We also evaluate using other hyperparameter settings, as well as a random forest classifier for $g$ instead, in Appendix B.2, where we found similar results.
@@
0:  Image $x\in\mathbb{R}^{C\times H\times W}$ with mask $m\in\{0,1\}^{H\times W}$, first convolutional layer of pretrained CNN $f_{1}$, simple classifier $g$.
@@
5:  $h_{\mathrm{obj}}:=h[:,m\neq 0]$
6:  $h_{\mathrm{bdry}}:=h[:,m^{\prime}\neq 0]$
7:  Train $g$ to classify between activations in $h_{\mathrm{obj}}$ vs. $h_{\mathrm{bdry}}$.
8:  return  Training classification accuracy of $g$.
@@ 3.2.1 Style Transfer for Controlled Textural Contrast
Similar to Sec. 2.3, in studying the effects of textural separability on SAM’s performance, we also wish to perform experiments on well-controlled synthetic images in order to more carefully disentangle the effects of textural separability shape on performance from the effects due to object shape, before evaluating on real data (in the following section). Here, we will do so using neural style transfer [15] to precisely modify the textural contrast of real objects/images. This will involve (1) modifying objects’ shape without altering their background using an inpainting model, followed by (2) applying neural style transfer (NST) to the composite image to adjust the textural contrast of the object with the surrounding image.
We begin with images sampled from the VOC2012 dataset [13] accompanied by objects (instance masks), according to certain criteria. Namely, we pick objects which take up between $5\%$ and $25\%$ of the entire image’s area (to have large enough objects while still maintaining sufficient background), sampling 484 objects total (with accompanying background) for study. For each object, we use a Stable Diffusion-based inpainting model (details in Appendix C.2) to remove the object from its corresponding image and fill in it’s mask area with the background. This creates a pair of a background image and a separate object (mask), allowing us to have careful control of modifying the object (and its mask accordingly) before placing it back into the inpainted background. For a given object, we then create three types of composite images with varying degrees of changes to the object’s shape (without modifying the background):
@@
    Altered: the object undergoes major non-affine geometric transformations (see App. C.2), resulting in shape modification and its texture being squeezed or expanded.
@@
    Mixed: this variant uses pixels from the altered image where the original and altered masks overlap, and pixels from the original object where they do not overlap.
After one of these three composite image types is created, we use neural style transfer (NST) to apply a texture to the image (similar to experiments in [16]) randomly sampled from Colored Brodatz [1], via an implementation of the NST model of Gatys et al. [15] (details in Appendix C.2). We illustrate an example of this entire procedure in Fig. 6.
Figure 6: Example images from each component of the synthetic textural separability dataset creation pipeline.
We adjust textural separability by performing the style transfer at eight different degrees of severity using different settings for the weighting hyperparameters that control the balance between content preservation and style transfer (details in Appendix C.2), with example styled images shown in Fig. 7 (more in Appendix. A.3). In Fig. 8 left, we validate this scheme by showing that indeed, steadily increasing the NST intensity results in decreasing textural separability.
Figure 7: Using style transfer with different intensities to adjust the textural separability of an object (highlighted in red), on controlled (upper row) and altered (lower row) versions of the object.
@@
In Fig. 8 right, we show how SAM’s segmentation performance changes with respect to style transfer intensity on the test set, for each of the three types of object transformations. We first see that SAM’s performance decreases with lower textural separability/higher NST intensity. Second, there is an interesting pattern of clear gaps in segmentation performance between the three object transformation types. SAM performs best on the altered objects/foregrounds, which have both distorted shape and texture, compared to the un-altered/controlled objects, with the mixed objects in-between. This further demonstrates the effect of textural contrast on segmentation performance, as the altered objects have additional textural contrast to their backgrounds due to the transformations which were applied to their shape, boundary and texture.
@@ 3.2.2 Experiments on Real Data
We will now evaluate datasets of real images; iShape (Sec. 2.3.2) and Plittersdorf [18], both of which were used in the original SAM paper [27], and possess objects with a wide range of textural separability from their surroundings. Plittersdorf consists of camera trap video frames of wild deer recorded in a wildlife park. These frames often have low contrast objects due to frequent low-light conditions, making it a useful dataset for this analysis. Example images and objects from both datasets are shown in Fig. 9.
In Fig. 10, we show how the textural separability (Algorithm 3) of these objects relates to SAM’s segmentation performance on them, with correlation results shown in Table 3. Overall, across all datasets we see a fairly strong correlation between textural separability and segmentation performance (average correlation of $\tau=0.49$ and $\rho=0.66$), especially considering the variety of objects and backgrounds, with objects with low separability resulting in especially poor segmentation ($\mathrm{IoU}\sim 0.3$). In these cases, SAM was confused by objects close to the object of interest which also have similar textures (see e.g., Fig. 10, right).
@@
Table 3: Rank correlation between SAM prediction IoU and object textural separability, on iShape and Plittersdorf.
@@
Figure 10: SAM prediction IoU vs. object textural separability (Algorithm 3), for iShape (left) and Plittersdorf (right).
Discussion
Understanding and Interpreting SAM’s Failure Modes.
@@ Discussion
First, we see that SAM’s performance issues on cases of high tree-likeness (Sec. 2.3) were primarily due to over-segmentation, rather than under-segmentation; the model would frequently generate large false-positive regions in the empty space of densely packed, thin tree-like structures (e.g., Figs. 3, right and 1, left). This could not simply be due to image resolution/downsampling blurring the thin objects, as for structures which are also thin but more sparsely packed, the over-segmentation is reduced (see e.g., Fig. 3, center rightmost). We therefore hypothesize that the root of SAM’s failure on these objects may lie in the highly repetitive yet irregular, dense patterns inherent to tree-like structures in particular: SAM confuses these patterns of shape as the texture of a more regularly-shaped object.
@@
In the cases of both tree-like and low-textural contrast objects, SAM either has trouble with correctly delineating an object’s shape when faced with a “confusing” textural cue, be it either a dense, irregular shape which appears to instead be a texture, or an object texture that is similar to its surroundings. One explanation for this is that SAM’s training set SA-1B [27] possessed few objects with these qualities. This seems likely for tree-like objects, given the typically low concavity of objects in SA-1B (Fig. 6, right in [27]), which were captured in photographic contexts where such objects are uncommon. Low-textural contrast objects are similarly uncommon in most photographs, and furthermore, segmenting objects is generally more challenging if they are harder to pick out from their surroundings.
Applications of Our Findings and Metrics.
A simple application of our findings and proposed metrics for the tree-likeness and textural separability of objects would be to predict if SAM is expected to perform underwhelmingly on new data, according to if the objects’ measured tree-likeness is high or the textural separability is low. Similarly, the general diversity of objects in a segmentation foundation model’s training/fine-tuning dataset could be measured according to the dataset’s distribution of these metrics, which could inform whether additional training data is needed to be acquired/annotated to develop a better generalist model.
Limitations and Future Work.
For our findings of the correlations between the object characteristics measured by our metrics and SAM’s performance in segmenting such objects, we attempted to minimize the influence of other confounding factors which could effect segmentation performance by first establishing a baseline with our carefully-controlled synthetic data experiments. We then chose real evaluation datasets which covered a wide range of object types according to these metrics in order to gain a better “signal-to-noise ratio” for the studied trends. However, it is impossible to mitigate all potential confounding factors which could affect performance in real data, which is why our correlation findings are still noticeable for the real datasets, yet not quite as tight as on the synthetic data.
We do not evaluate the 2D version of the recently released SAM 2 model due to it likely resulting in similar behavior as SAM 1, along with evidence that oracle predictions of SAM and SAM 2 are typically similar in performance to within $\lesssim 0.1$ IoU [12].
Another future direction would be to explore the specific component(s) of SAM which can be most attributed to these failures (in a mechanistic sense), which could guide further model development. We deem it likely that the susceptible component of SAM is the image encoder, simply because it is the main backbone of the model which extracts features to be used by the lightweight mask decoder, but obtaining a more fine-grained answer would require a careful treatment.
Conclusion
In this paper, we quantitatively modeled how the segmentation performance of SAM relates to certain measurable object characteristics: tree-likeness and textural separability. We find SAM’s performance to typically be noticeably correlated with these factors, showing the need for further work in understanding and potentially mitigating such “failure modes” of vision foundation models.
@@ Appendix C Additional Experimental Details
C.1 SAM Prompting Strategies
In addition to the SAM prompting details presented in Sec. 1, all real and style-transferred (NST) images are prompted with (1) a tight bounding-box and (2) several positive and/or negative prompts randomly sampled from either the foreground or background of the object mask, respectively, with respective counts $n_{pos}$ and $n_{neg}$. For iShape, we use $n_{pos}=n_{neg}=5$; for DIS, we use $n_{pos}=5$ and $n_{neg}=10$ due to the complexity of the objects. For Plittersdorf and NST images, we simply use $n_{pos}=0$ and $n_{neg}=2$, due to the objects typically possessing simple shapes.

### v2

@@ 5 The Challenge of Textural Separability
In the previous section, we demonstrated that SFMs struggle with segmenting non-conventional shapes, in particular, dense tree-like structures, which we hypothesize is due to the model confusing the dense structure as the texture of a non-treelike, more regular shape, rather than a shape itself (see e.g. Fig. 1 left). Similar to this behavior is that even for objects with simpler shapes, these SFMs can still be confused if the object’s texture is even somewhat similar to its surroundings, which we will study in this section.
@@ 5.1 Measuring Textural Separability
We will define the textural contrast or separability between some object mask and its surroundings by how easily their textures can be distinguished from one another. Motivated by findings that early layers of classification-pretrained CNNs primarily capture low-level features involving edges and textures [47], we characterize an image’s textures using the first convolutional layer of a ResNet-18 [18] pretrained on ImageNet [9]. Denoted by $f_{1}:\mathbb{R}^{C\times H\times W}\rightarrow\mathbb{R}^{C^{\prime}\times H^{\prime}\times W^{\prime}}$, this outputs a textural feature map corresponding to all $7\times 7$ windows in the image.
We then measure the textural separability of an object according to if a simple classifier $g$ can be trained to discriminate between (a) the activations for the pixels of the object foreground and (b) the activations right outside of the object boundary. We define this explicitly in Algorithm 3, where $\mathrm{Dilate}$ and $\mathrm{Disk}$ refer to scikit-image.morphology functions [43]. We use simple logistic regression for $g$ with inverse regularization parameter $C=2$ [32]. We also evaluate using other hyperparameter settings, as well as a random forest classifier for $g$ instead, in Appendix B.2, where we found similar results.
@@
0:  Image $x$ with binary mask $m$, first convolutional layer of pre-trained CNN $f_{1}$, simple classifier $g$.
@@
5:  $h_{\mathrm{obj}}:=h[:,m\neq 0]$, $h_{\mathrm{bdry}}:=h[:,m^{\prime}\neq 0]$
6:  Train $g$ to classify between activations in $h_{\mathrm{obj}}$ vs. $h_{\mathrm{bdry}}$.
7:  return  Training classification accuracy of $g$.
@@ 5.2.1 Style Transfer for Controlled Textural Contrast
To isolate the effects of textural separability from object shape, we first conduct controlled experiments using neural style transfer (NST) [14], before evaluating on real data in the next section. This allows us to systematically adjust an object’s textural contrast with its background while keeping other factors controlled.
We begin with all objects sampled from VOC2012 [12] that occupy 5–25% of their image area, to ensure sufficient background context (484 objects total). Using a Stable Diffusion-based inpainting model (App. C.2), we remove each object and reconstruct its background, creating pairs of separate object-only and background-only images. This setup allows precise modifications before reinserting objects into their original scenes. For each object, we then create three types of composite images with varying degrees of changes to the object’s shape (without modifying the background):
@@
    Altered: the object undergoes significant non-affine geometric transformations, modifying its shape and distorting its texture (App. C.2).
@@
    Mixed: a hybrid of the original and altered object, using altered pixels where the masks overlap and original pixels elsewhere.
After creating these image variants, we apply NST (similar to [15]) using textures randomly sampled from Colored Brodatz [1] with the Gatys et al. NST model [14] (details in App. C.2). An example of this entire procedure is in Fig. 6.
Figure 6: Example from the NST image creation pipeline.
To systematically vary textural separability, we apply NST at eight different intensity levels, adjusting the balance between content preservation and style transfer (details in App. C.2). Fig. 7 provides examples of this, and Fig. 8 (left) confirms that increasing NST intensity steadily reduces textural separability as intended.
Figure 7: Using NST with different intensities to adjust the textural separability of an object (highlighted in red), on controlled (upper row) and altered (lower row) versions of the object.
@@
In Fig. 8 right, we analyze how SAM’s performance varies with NST intensity across the three object transformation types. First, as textural separability decreases (higher NST intensity), SAM’s performance declines, confirming the impact of textural contrast on segmentation accuracy. Second, we observe clear performance gaps between transformation types: SAM performs best on altered objects, which undergo both shape and texture distortion, followed closely by mixed objects, and then performs worst on controlled objects. This suggests that additional textural contrast—introduced by shape, boundary, and texture transformations—enhances segmentation performance, further reinforcing the role of textural cues in SFM behavior.
@@ 5.2.2 Experiments on Real Data
We will now evaluate datasets of real images; iShape (Sec. 4.3.2) and Plittersdorf [17], both of which were used in the original SAM paper [26], and possess objects with a wide range of textural separability from their surroundings. Plittersdorf consists of camera trap video frames of wild deer recorded in a wildlife park. These frames sometimes have low-contrast objects due to frequent low-light conditions, making it a useful dataset to analyze. Example images and objects from both datasets are shown in App. A.4.
In Fig. 9, we show how the textural separability (Algorithm 3) of these objects relates to SAM’s performance on them, with correlation results shown for all three SFMs in Table 3. Overall, across all datasets, we see a fairly strong correlation between textural separability and segmentation IoU (average corr. of $\tau=0.47$ and $\rho=0.62$), especially considering the variety of objects and backgrounds, with objects with low separability typically being noticeably harder to segment. In these cases, the SFMs were confused by objects close to the object of interest which also have similar textures.
@@
Table 3: Rank correlation between SFM IoU and object textural separability, on iShape and Plittersdorf.
@@
Figure 9: Segmentation IoU vs. object textural separability, for iShape (left) shown for SAM, and Plittersdorf (right) for all SFMs.
6 Robustness to Fine-tuning
Despite being designed for thin, branching structures, HQ-SAM exhibits performance drops on high tree-likeness objects, similar to SAM and SAM 2 (Table 2, Fig. 5). This suggests that the difficulty SFMs face with such objects persists even after adaptation, a hypothesis we explore here.
To assess whether fine-tuning could improve SAM’s segmentation of high tree-likeness objects, we selected some $k$ examples from DIS-train using different ranking strategies: Oracle (direct IoU-based selection), Rand (random selection), DoGD / CPR (selecting examples based on high tree-likeness), and Sum (a combination of DoGD and CPR). We fine-tuned only SAM’s mask decoder on these, keeping the image encoder frozen to preserve its generalist capabilities (as in [24]); full details are in App. C.3.
Results (Fig. 10) show that fine-tuning on high tree-likeness examples (DoGD, CPR, or Sum) led to slightly worse validation performance than random selection, likely due to overfitting on these atypical examples, harming generalization. Even though fine-tuning improved overall average IoU, the correlation between segmentation performance and tree-likeness remained—both on unseen validation data (middle plots, Fig. 10) and even on the fine-tuned examples themselves (lower plots). These findings indicate that fine-tuning alone is not a viable solution to overcoming SFMs’ performance drop on highly tree-like objects.
Figure 10: Top: Performance (IoU) of fine-tuned SAM on DIS val vs. number of fine-tuning examples $k$ sampled from DIS train, using different selection strategies. Lower rows: IoU vs. tree-likeness for each $k=1024$ fine-tuned model on DIS val (middle row) and on the fine-tuning set (bottom row).
7 Discussion
@@
Understanding SFM Failure Modes. We find that SFMs struggle primarily with over-segmentation rather than under-segmentation in high tree-likeness cases (Sec. 4.3). The models often generate large false-positive regions within densely packed, thin structures (e.g., Figs. 3, right; 1, left). This cannot be explained by resolution limitations alone, as sparsely packed thin structures exhibit reduced over-segmentation (Fig. 3, center rightmost). Instead, we hypothesize that SFMs misinterpret dense, irregular patterns as texture, confusing these structures with the appearance of a more uniformly shaped object.
@@
In the cases of both tree-like and low-textural contrast objects, SFMs struggle to delineate object boundaries when faced with misleading textural cues, be it either a dense, irregular shape that appears to instead be a texture (the former case), or an object texture that is similar to its surroundings (the latter). This suggests that SAM’s training set (SA-1B) contained few such objects, as tree-like structures and low-contrast objects are uncommon in typical photographic datasets [26] (for the former, see e.g., the typically low object concavity of SA-1B in Fig. 6, right in [26]). Since HQ-SAM and SAM 2 were derived from SAM/its training set, their similar behavior is unsurprising.
On HQ-SAM and Fine-tuning. Despite being designed and adapted for thin, intricate structures [24], HQ-SAM still exhibits performance drops correlated with tree-likeness (Tables 1, 2), albeit with slight improvements over SAM and SAM 2 (Fig. 5, left). This may stem from HQ-SAM’s adapter-based approach, which leaves SAM’s image encoder and mask decoder frozen, preserving SAM’s limitations. Our fine-tuning experiments (Sec. 6) further suggest that this issue is not easily mitigated, indicating that these structures pose a fundamental segmentation challenge.
Limitations and Future Work. To ensure our findings were robust, we minimized confounding factors by first establishing trends in controlled synthetic experiments before validating them on real datasets spanning diverse object types. However, real-world segmentation is inherently complex, making it impossible to eliminate all confounding variables—which may explain why observed correlations were stronger in synthetic data than in real datasets.
A key future direction is to identify the specific model components responsible for these failures. SAM’s image encoder certainly plays a central role, as it serves as the primary feature extractor for the lightweight mask decoder, but a more precise mechanistic analysis could guide architectural improvements to better handle complex structures.
Conclusion. In this paper, we quantitatively modeled how SFM segmentation performance relates to tree-likeness and textural separability. Our findings reveal a consistent correlation between these object characteristics and segmentation accuracy, highlighting the need for further research to better understand and mitigate these failure modes—especially since even targeted fine-tuning fails to fully resolve them.
@@ A.3 Neural Style Transfer (NST) Generated Images
A.4 Example Images from Plittersdorf and iShape
We show example images from Plittersdorf and iShape in Fig. 13.
@@ Appendix C Additional Experimental Details
C.1 SFM Prompting Strategies
In addition to the SFM prompting details presented in Sec. 3, all real and style-transferred (NST) images are prompted with (1) a tight bounding-box and (2) several positive and/or negative prompts randomly sampled from either the foreground or background of the object mask, respectively, with respective counts $n_{pos}$ and $n_{neg}$. For iShape, we use $n_{pos}=n_{neg}=5$; for DIS, we use $n_{pos}=5$ and $n_{neg}=10$ due to the complexity of the objects. For Plittersdorf and NST images, we simply use $n_{pos}=0$ and $n_{neg}=2$, due to the objects typically possessing simple shapes.
@@ C.2 Neural Style Transfer Experiments
C.3 Fine-Tuning Experiment Details
Here we provide all details for the SAM fine-tuning experiments of Section 6. The loss function that we use is the sum of the Dice and cross-entropy losses (as implemented by MONAI [5]), optimized via AdamW [29] with a learning rate of $10^{-5}$ and a weight decay strength of $0.1$. We use a batch size of $16$ and fine-tune for $40$ epochs on the $k$ (image, mask) examples which were sampled from DIS-train using the given selection strategy (see Section 6). All parameters of SAM’s mask decoder are fine-tuned, while the image encoder is frozen.
@@ D.1 Correlation Between Metrics
D.2 Effect of Object Line Thickness on Performance
In Section 4.1, we hypothesized that the noticeable SAM segmentation performance difference between retinal vessel images and satellite road images, two types of objects with branching structures, is due to the former objects typically having more dense and irregularly-spaced branching structures than the latter. However, it is important to consider the possibility that this is simply due to retinal vessels having thinner lines than road objects, which would help evaluate if the proposed tree-likeness metrics (CPR and DoGD) are actually characterizing the true underlying features that drive performance.
To test this, we performed the same synthetic data experiments as in Section 4.3.1, but applied transformations which thickened object lines, to see how this affected the overall relationship between measured object tree-likeness and segmentation performance. To do so, between steps 2 and 3 of the procedure used to generate these images (App. A.2), we:
  1. 1.
    Converted the object mask $m_{c}$ to its skeleton, effectively resulting in a mask consisting of lines with 1-pixel widths.
  2. 2.
    Apply a dilation transformation with a radius of 4 pixels to this skeleton, resulting in each path/line structure in the mask being strictly 9 pixels wide.
We show the results of this procedure (how it affects the relationship between segmentation IoU and tree-likeness) in Fig. 15. We see that the result of IoU and object tree-likeness being correlated is mostly invariant to the object line thickening transformation. However, while the DoGD results are generally maintained because DoGD quantifies the deviation of filling-rate at different scales, we see that CPR experiences some degradation in its correlation with IoU for certain retinal vessel components. This is possibly due to the highly dense tree structures, originally thin with varying widths, forced to be similarly thickened and then possibly merged with each other, thereby reducing the overall CPR of the object. This may be a limitaiton of CPR in quantifying the tree or net-like geometric structure. Fortunately, such special cases where all object pixels fall around a fixed range of object skeletons are rare in realistic scenarios.
Figure 15: SAM segmentation IoU vs. object tree-likeness (CPR, left; DoGD, right) on the synthetic dataset with dilated object lines.